I have a question that I would like answered please....
Solve for x:
2(x-3) = 8(4-x) =38
Thank you for your help in answering this problem.!!! :)
nevrstndstil@yahoo.com

Answer by vertciel(183)

Hi there,
I recategorised your question. This will help future users and you, because your question will be answered more quickly.
I assume your equal sign in the middle means the minus sign: 2(x-3) - 8(4-x) = 38.
2(x - 3) - 8(4 - x) = 38
2x - 6 - 8(4 - x) = 38
2x - 6 - 32 + 8x = 38
Now just organise the terms and isolate x. Have fun!